OpenStudy (anonymous):

Subtract. Express your answer in simplest terms. 8 4/5 - 5 1/10

OpenStudy (anonymous):

really

OpenStudy (anonymous):

well im in 5th grade

OpenStudy (kidrah69):

Start by turning the fractions into improper fractions

OpenStudy (kidrah69):

\[\huge~\rm~Unit \times denominator+numerator \]

OpenStudy (anonymous):

oh i found the answer thank you

OpenStudy (kidrah69):

Just double check your work \[\huge~\rm~\bf~\frac{ 44 }{ 5 }-\frac{ 51 }{10 }\] We need to find a common denominator. which is 10 \[\huge~\rm~\bf~\frac{ (2)44 }{ (2)5 }-\frac{ 51 }{10 }\] \[\huge~\rm~\bf~\frac{ 88-51 }{ 10}=\frac{ 37 }{ 10 }\]

OpenStudy (anonymous):

sorry